A famous IT consulting firm BracKenzie and Co, henceforth called the Firm, has a consulting project implementing ERP at a large telecom firm, henceforth called the Client. The project manager, Mr. Grisham, is in the process of submitting a staffing proposal to his client for the project. The labor requirement is estimated to follow a normal distribution with mean 3000 person-hours and a standard deviation of 400 person-hours.

Because the Client is large and influential, the Client pays the Firm for exactly as many hours as are required to complete the project. It is unlikely however, that unutilized time of the consultants will be used up on other projects on short notice, and therefore the Firm will continue to pay $175 per hour to its project staff, per hour without any revenue.

On the other hand, since the project is high-profile, if the work requirements exceed the budgeted person-hours, the Client will hold the Firm accountable for project completion. The Firm will then have to pay $250 per person-hour to hire contract consultants on short notice to make-up for the remaining work requirements.

  1. Mr. Grisham has asked you for advice on how many person-hours the Firm should allocate to this project with the Client?

The Firm Should Allocate                                     Person-Hours

  1. What would the hourly rate for the contract consultants have to be (as opposed to the current rate of $250), in order for you to advise Mr. Grisham to allocate exactly 3000 person-hours to the project?

  

            $                                  per Person-Hour

  1. Given your answer in part (a), what is the expected number of contract consultant hours to be utilized (at the hourly rate of $250)?

Expected number of contract consultant hours is                                    
Answer #1
  1. Mr Grisham has asked you for advice on how many person-hours the Firm should allocate to this project with the Client?

The Firm Should Allocate Person-Hours:

Mean

Standard Deviation

Min Person per Hour

Max Person per Hour

3000

400

2600

3400

There is a possibility of the requirement of 3400 people per hour in certain circumstances, considering the mean 3000 people per hour and Standard deviation 400 Person per hour.

In order to avoid shortage during the project, it is advisable to have 3400 people per hour.

  1. What would the hourly rate for the contract consultants have to be (as opposed to the current rate of $250), in order for you to advise Mr Grisham to allocate exactly 3000 person-hours to the project?

Given: The Firm will continue to pay $175 per hour to its project staff, per hour without any revenue. But in case of the unexpected situation, the firm may be forced to pay $250 if the project is not completed on time.

I would recommend paying $175 per hour even though the client couldn’t complete within the given time frame. So that the client will try to complete on time.

$per Person-Hour = 3000*175 = $525000 for project or $175 per person hour

  1. Given your answer in part (a), what is the expected number of contract consultant hours to be utilized (at the hourly rate of $250)?

Expected number of contract consultant hours = 3400 * 250 = $850,000